OpenStudy (australopithecus):

3. A reservoir contains 600 litres of pure water. Brine (salty water) that contains 0.2 kg/L of salt is added at a rate of 2 L/min. Brine from a second source with 0.05 kg/L of salt is added at a rate of 3 L/min. Assume that the reservoir is instantaneously well-mixed. The reservoir is drained at a rate of 5 L/min. Let Q(t) be the amount of salt (in kg) at time t (in min). (a) Set up and solve the differential equation for Q(t).

OpenStudy (australopithecus):

so I know the general formula is Rin - Rout = dy/dt

OpenStudy (australopithecus):

but yeah I'm stuck

OpenStudy (dumbcow):

well you know the volume remains constant at 600 combined sources flowing in is 5L/min and flows out at 5L/min Rate of salt going in is (.2*2 = .4) + (.05*3 = .15) --> 0.55 kg/min Rate of salt going out is 5*concentration at time t --> 5*(Q/600) = Q/120 kg/min thus \[\frac{dQ}{dt} = 0.55 - \frac{Q}{120}\]

OpenStudy (australopithecus):

thanks, that makes sense :) so the Rin is normally a constant and Rout is always the outflow*concentration at time t/volume

OpenStudy (dumbcow):

yep

OpenStudy (australopithecus):

if there was an initial concentration in the chamber like 0.2kg/L how would I represent that in the equation

OpenStudy (dumbcow):

to solve this DE you need to find the integrating factor ...hmm, i don;t think it would change the rate equation just the function Q(t) because you would apply different initial conditions

OpenStudy (australopithecus):

I know how to solve these problems I just need help setting up the equations

OpenStudy (australopithecus):

so it would be the same if it was 0.2kg/L in the tank instead of pure water initially?

OpenStudy (dumbcow):

yes, because dq/dt is rate that salt is changing over time concentration is ratio of salt per volume so its an initial condition and does not affect the rate the salt changes once you start pumping in/out water

OpenStudy (dumbcow):

obviously the value of Q in the equation will be different but you would still use same rate equation

OpenStudy (australopithecus):

ok thanks

OpenStudy (dumbcow):

no prob

OpenStudy (australopithecus):

I got the equation Q = 66e^(x/120) - 66, does this seem reasonable? I dont think I trust it